# Personalized Wireless Federated Learning for Large Language Models

### v1

@@ IV-A Current Research Progress
There have been several joint research efforts on wireless FL and LLMs. The authors in [3] combined split learning and FL to pretrain the BERT model. Reference [4] proposed a federated instruction learning called Shepherd, which utilizes LoRA for fine-tuning LLMs through instruction data. Similarly, the authors in [8] presented a low-parameter federated learning based on LoRA for task fine-tuning of LLMs. Reference [10] addressed co-tuning and off Site-tuning of LLMs through a comprehensive FL open-source framework called Fate-LLM. However, all of these studies aim to train a unified LLM in a distributed manner, overlooking device variations, user preferences, and characteristics. As a result, they fail to achieve device-adaptive and user-centric LLMs.
@@ IV-B1 Personalized User Data
PFL allows personalized learning from various user data, which can be beneficial in fine-tuning LLMs on non-identically and independently distributed (non-IID) data. PFL can learn a personalized LLM for each client that is tailored to its own data distribution, enabling the LLM to learn from a wider range of personalized contexts, features and patterns. This can lead to improved understanding and generation capabilities of LLMs for user private data.
@@ IV-B2 Customized Local Model
By allowing each client to have its own personalized LLM, PFL enables customization of the local tuning process to each device’s preference and constraints. This can lead to better suitability to personal computational, storage, and communication resources and improved model performance for all clients.
@@ IV-B3 Specific Communication Process
Therefore, PFL provides flexibility in balancing the trade-off between learning shared knowledge and personalized knowledge for LLMs.
IV-C Personalized Federated Instruction Tuning
@@
We propose a PFIT method based on RL, in which each client has personalized requirements for helpfulness and safety of the LLM. Helpfulness emphasizes the quality and accuracy of generated content, such as grammatical correctness, logical coherence, and relevance of the responses. Safety, on the other hand, emphasizes the legality and ethicality of the generated content, such as the absence of sensitive or harmful information. For example, in Fig. 2, the LLM with high helpfulness would provide answers to user questions regarding employee information, including sensitive details. However, the LLM with high safety levels would refrain from answering sensitive employee information, prioritizing user privacy and data protection. To achieve PFIT, we introduce three key innovations:
  • Double Reward Model: We define two reward models to evaluate the helpfulness and safety of local instruction responses. The quality reward for different clients’ outputs is obtained by linearly weighting these two reward models, enabling personalized instruction fine-tuning of the LLM for different clients’ requirements.
  • Personalized Reward Function: We design a personalized reward function that includes both the quality reward from two reward models and a negative regularization reward for global knowledge. The regularization reward is based on the Euclidean distance between local model parameters and global model parameters, promoting the knowledge sharing among clients in the PFL system.
  • Sparse Attention Update: To encourage lightweight devices to participate in PFL, we only train the last two layers of the LLM. Additionally, we adopt a sparse attention mechanism to further reduce the computation complexity and communication overhead in self-attention layers during the fine-tuning phase of the LLM.
The specific workflow of the PFIT is as follows:
Step 1: Initialize the pre-trained LLM as the global model on the server and freeze the earlier parts of the LLM.
Step 2: Each client uses the global LLM as the initial local LLM, sets a personalized reward function (In the red dashed box of Fig. 2), and selects its own instruction data for local learning.
Step 3: Based on the current local LLM, the client calculates the regularization reward related to the global LLM, evaluates the helpfulness and safety of responses, computes the quality reward for instructions, and then updates the unfrozen part of the local LLM using the PPO algorithm [11] according to the personalize reward function, including both quality reward and regularization reward.
Step 4: The server aggregates the sparse tunable layers (the unfrozen parts of LLMs) from all clients, obtains the updated global LLM, and sends the global LLM (the unfrozen part) to all clients.
Step 5: Steps 3-4 are repeated until the convergence criteria of the PFL system are met.
Figure 2: The illustration of the proposed PFIT based on RL.
IV-D Personalized Federated Task Tuning
We propose a PFTT method based on PEFT, in which a set of clients have different task goals, and each client possesses a set of non-IID task data. For instance, in Fig. 3, different clients are tasked with film classification, but the distribution of labelled data varies across different clients. Some clients may have a higher proportion of science fiction and realistic film labels, while others may have a larger number of comedy and tragedy film labels. As a result, the LLM will excel at providing personalized movie classifications based on the local distribution of labelled data available to it. To achieve PFTT, we introduce three key innovations:
  • Universal Adapter: We incorporate adapters into the global pre-trained LLM to share task knowledge among different clients.
  • Local LoRA: We introduce LoRA in the local LLMs to enable personalized local knowledge learning. The size of LoRA can be adjusted based on the data volume or computational resources available on each client.
@@ IV-D Personalized Federated Task Tuning
  • Partial Aggregation: During global aggregation, only the adapter parameters are aggregated for global knowledge sharing, while the LoRA parameters are not aggregated for maintaining personalization. This approach enables personalized task tuning in heterogeneous devices.
The specific workflow of the PFTT is as follows:
@@
Step 1: Initialize the pre-trained LLM as the global LLM on the server and insert adapters to the LLM.
@@
Step 2: Each client uses the global LLM as the initial local LLM and designs local LoRA parameters based on the data volume or computational resource of the local LLM.
@@
Step 3: Based on the current global LLM and local LoRA parameters, the client fine-tunes the LLM using local task data and updates the adapter and LoRA parameters over the wireless networks.
@@
Step 4: The server aggregates the adapter parameters from all clients, obtains the updated global adapter parameters, and sends them to the clients.
Step 5: Steps 3-4 are repeated until the convergence criteria of the system are met.
Through PFL, both fine-tuning methods can achieve better performance for personalized LLMs.
Figure 3: The illustration of the proposed PFTT based on PEFT.
@@ V Simulation Results
This section provides two experiments for demonstrating the effectiveness of the proposed PFIT and PFTT.
V-A Problem Formulation
@@
We consider a PFL system for LLMs. Suppose the system consists of four clients and one server, where each client possesses different local data and model preferences. The server has sufficient resources. The PFL system aims to achieve personalized fine-tuning of the local LLMs for all clients in wireless networks with Rayleigh channel, the communication round is set to 40 and the SNR is set to 5dB.
V-B Simulation Settings
V-B1 Settings for PFIT
We first employ the Alpaca dataset [12] to evaluate the validity of the proposed PFIT scheme. Secondly, we use GPT-2 [13] as the local LLM with 40% sparse attention and adopt PPO as the local RL algorithm. During the fine-tuning process, we sample instructions from the dataset to generate the corresponding responses from GPT-2. Then, we incorporate the “instruction+response” into two reward models, with one model assessing the helpfulness score of the response and the other evaluating its safety score. We then utilize the reward score (i.e., helpfulness score plus safety score) and communication cost (i.e., the size of parameters to be aggregated) per round as evaluation metrics.
V-B2 Settings for PFTT
We first employ the AG’s news corpus [14] as the evaluation dataset for PFTT. Additionally, we adopt a Dirichlet distribution to facilitate a non-IID data partition among clients. Next, we utilize RoBERTa [15] as the local LLM, which is an improved BERT model pre-trained on a substantial corpus of English data in a self-supervised manner. We use 12 universal adapters for each client to exchange information among different clients. Subsequently, each client incorporates 10-12 local LoRAs, based on their local resources, to achieve local model personalization. Lastly, since the local LLM is responsible for sentence classification, we use the classification accuracy and communication (i.e., communication cost divided by transmission rate) delay per round as evaluation metrics.
V-C Evaluation Results
Fig. 4 presents the evaluation results for PFIT and its contenders, where SFL means to a fine-tuning method that uses only a single reward model (helpfulness metric) and incorporates 20% sparse attention. PFL, on the other hand, represents personalized fine-tuning without the sparse attention. Shepherd is an FL method that employs LoRA for instruction fine-tuning [4]. We can see that PFIT enables the local model to obtain the highest rewards than SFL and Shepherd with the single reward model. Moreover, compared to PFL that does not use sparse attention, PFIT reduces communication overhead by 20%. Shepherd, utilizing LoRA for instruction fine-tuning, results in the lowest communication overhead. However, this approach also affects the LLM’s performance, resulting in lower rewards compared to PFIT.
Figure 4: Reward and communication cost for PFIT and its contenders.
@@ V-C Evaluation Results
Fig. 5 presents the evaluation results for PFTT and its benchmarks. In vanilla FL [1], the parameters of both adapters and LoRAs all need to be uploaded. FedBert [3] is a federal split learning method, and FedLora [8] is a federated fine-tuning method that exclusively incorporates LoRA. The results indicate that PFTT achieves the highest accuracy, which highlights the effectiveness of the LoRA-based personalized structure. Similarly, due to the fact that PFTT only requires the transmission of a part of fine-tuned parameters (universal adapters), it incurs the minimum communication overhead compared to other methods.
Figure 5: Accuracy and communication cost for PFTT and its contenders.
@@ VI-1 Wireless Aggregation and Divergence
In PFL, multiple participants collaborate to train a shared LLM. However, due to the possibility of signal quality fluctuations in wireless networks, mobile devices may experience communication interruptions and data loss. In the aggregation, these instabilities can lead to model divergence, where participants have different contributions to model updates. Addressing the challenges of model aggregation and divergence in wireless networks requires asynchronous model aggregation strategies and fair client selection mechanisms to ensure the model effectively incorporates contributions from all participants while balancing the differences to ensure the reliability of model updates.
@@ VI-2 Personalization and Overfitting
Personalization is one of the core objectives in PFL, aiming to customize the shared LLM to meet specific needs of each client. However, the introduction of personalization can potentially lead to overfitting problems. If the personalized requirements are overly detailed, the LLM may overfit to the data of specific clients, resulting in degraded performance on other clients or tasks. Resolving the challenges of personalization and overfitting necessitates appropriate regularization and control during the fine-tuning process to strike a balance between the degree of personalization and the LLM’s generalization.
@@ VI-3 Communication Efficiency and Model Accuracy
PFL involves communication and collaboration among multiple participants. Communication overhead can be a significant challenge, particularly in scenarios with LLMs and numerous participants. Frequent communication can increase communication latency and resource consumption. Moreover, the unreliability or instability of communication can result in the loss or delay of model updates, which can have a direct impact on the accuracy and performance of the LLM. Addressing the issue requires designing efficient communication protocols and strategies to reduce communication overhead while ensuring reliable transmission of data and model parameters.
@@ VII Conclusion
In this paper, we have first summarized three learning stages of LLMs and discussed potential solutions for combining FL with LLMs. Next, we have proposed two PFLs for different fine-tuning methods. Specifically, we introduced PFIT for instruction fine-tuning of LLMs based on client preferences. We then designed two reward models based on helpfulness and safety, and we used RLHF to fine-tune the LLM based on diverse combinations of reward models from clients. Furthermore, we proposed PFTT to fine-tune downstream classification tasks based on locally non-IID data. In PFTT, we used global adapters to enable information exchange between devices and incorporated locals LoRA to customize the local LLM. Finally, we carried out simulations to validate the effectiveness of these proposed methods.
@@
Feibo Jiang (jiangfb@hunnu.edu.cn) received Ph.D. degree from the Central South University, China. He is currently an Associate Professor at Hunan Normal University, China.
Li Dong (Dlj2017@hunnu.edu.cn) received Ph.D. degree from the Central South University, China. She is currently an Associate Professor at Hunan University of Technology and Business, China.
@@
Kezhi Wang (Kezhi.Wang@brunel.ac.uk) received Ph.D. degree from University of Warwick, U.K. in 2015. Currently he is a Senior Lecturer with the Department of Computer Science, Brunel University London, U.K.
Kun Yang (kunyang@essex.ac.uk) received his PhD from the Department of Electronic & Electrical Engineering of University College London (UCL), UK. He is currently a Chair Professor in the School of Computer Science & Electronic Engineering, University of Essex, UK. He is also an affiliated professor at UESTC, China.
Cunhua Pan (cpan@seu.edu.cn) received Ph.D. degrees from Southeast University, China, in 2015. He held a post-doctoral position at Queen Mary University of London, U.K., from 2016 and 2019. From 2019 to 2021, he was a Lecturer in the same university. From 2021, he is a full professor in Southeast University, China.
Dusit Niyato (dniyato@ntu.edu.sg) received the Ph.D. degree in electrical and computer engineering from the University of Manitoba in Canada in 2008. He is a professor in the College of Computing and Data Science, Nanyang Technological University, 639798 Singapore.

### v2

@@ IV-A Current Research Progress
There have been several joint research efforts on wireless FL and LLMs. Tian et al. [5] combined split learning and FL to pretrain the BERT model. Zhang et al. [6] proposed a federated instruction learning called Shepherd, which utilizes LoRA for fine-tuning LLMs through instruction data. Similarly, Jiang et al. [8] presented a low-parameter federated learning based on LoRA for task fine-tuning of LLMs. Fan et al. [3] addressed co-tuning and off-site-tuning of LLMs through a comprehensive FL open-source framework called Fate-LLM. However, all of these studies aim to train a unified LLM in a distributed manner, overlooking wireless environments, user preferences, and characteristics. As a result, they fail to achieve device-adaptive and user-centric LLMs.
@@ IV-B1 Personalized User Data
PFL allows personalized learning from various user data, which can be beneficial in fine-tuning LLMs on non-identically and independently distributed (non-IID) data. PFL can learn a personalized LLM for each client that is tailored to its own data distribution, enabling the LLM to learn from a wider range of personalized contexts, features and patterns. This can lead to improved understanding and generation capabilities of LLMs for private data.
@@ IV-B2 Customized Local Model
By allowing each client to have its own personalized LLM, PFL enables customization of the local tuning process to each device’s preference and constraints. This can lead to better suitability for personal computational, storage, and communication resources and improved model performance for all clients.
@@ IV-B3 Specific Communication Process
Therefore, PFL offers flexibility in balancing the trade-off between shared knowledge in distributed learning and personalized knowledge for LLMs.
IV-C Key Innovations of PWFF
@@
We propose a PWFF method, in which each client has personalized requirements for the helpfulness and harmlessness of the LLM. Helpfulness emphasizes the quality and accuracy of generated content, such as grammatical correctness, logical coherence, and relevance of the responses. Harmlessness, on the other hand, emphasizes the legality and ethicality of the generated content, such as the absence of sensitive or harmful information. For example, in Fig. 2, the LLM with high helpfulness would provide instructions on how to make bombs, even if doing so violates human laws and ethical values. However, the LLM with a high level of harmlessness would refuse to address such inquiries. To achieve PWFF, we introduce three key innovations in Fig. 2 as follows:
@@
  • Global Partial Aggregation: To encourage lightweight devices to participate in PFL, we adopt the adapter and LoRA to fine-tune the local LLM. During global aggregation, only the adapter parameters are aggregated for global knowledge sharing, while the LoRA parameters are not aggregated for maintaining personalization. This approach enables personalized fine-tuning and further reduces the energy consumption and communication overhead.
  • Personalized Reward Model: We define two reward models to evaluate the helpfulness and harmlessness of local instruction responses of LLMs. The total reward for different clients’ outputs is obtained by linearly weighting these reward models, enabling personalized fine-tuning of the LLM for different clients’ preferences.
  • Local Multi-objective Alignment: We design a personalized loss for RL (e.g., PPO algorithm [11]) that includes two reward models (to evaluate immediate rewards) and two critic models (to predict future rewards). The combination of these four components enables more accurate personalized value estimation for each client, thereby ensuring stable and effective policy updates for the actor model.
Figure 2: The innovations of the proposed PWFF.
IV-D The Detailed Workflow of PWFF
The detailed workflow of the PWFF framework is presented in Fig. 3, as outlined below:
Figure 3: The workflow of the proposed PWFF.
Step 1: Federated Instruction Tuning:
@@ IV-D The Detailed Workflow of PWFF
Step 1.1: Initialize the pre-trained LLM as the global LLM on the server and insert adapters into the LLM.
@@
Step 1.2: Each client uses the global LLM as the initial local LLM and designs local LoRA parameters based on the data volume or computational resource of the local LLM.
@@
Step 1.3: Based on the current global LLM and local adapter and LoRA parameters, the client fine-tunes the LLM using local instruction data and updates the adapter and LoRA parameters, and then the client transmits the adapter parameters over the wireless networks.
@@
Step 1.4: The server aggregates the adapter parameters from all clients, obtains the global adapter parameters, and sends them to the clients.
Step 1.5: Steps 1.3-1.4 are repeated until the convergence criteria of the system are met.
Step 2: Federated Reward Learning:
Step 2.1: A small fully-connected network is designed as a reward model, which is used to map outcomes generated by the LLM to a numerical reward signal. For example, outputs with high helpfulness or harmlessness correspond to high rewards, while outputs with low helpfulness or harmfulness correspond to low rewards.
Step 2.2: On the client, preference data regarding helpfulness and harmlessness are collected separately, enabling the reward model to understand the relative quality of each outcome from the LLM.
Step 2.3: Each client trains two reward models using supervised learning with preference data, one to predict the effectiveness reward and the other to predict the harmlessness reward. The parameters of the reward models are then transmitted over the wireless network to the server.
Step 2.4: The server aggregates all reward model parameters from the clients to obtain global reward models, and then sends them back to the clients.
Step 2.5: Steps 2.3-2.4 are repeated until the convergence criteria of the system are met.
Step 3: Federated Alignment Tuning:
Step 3.1: The working process of the LLM is modeled as a Markov Decision Process (MDP). Each client freezes its two reward models to predict the immediate reward for the current behavior of the LLM, while also designing two learnable Critic models based on these reward models to predict the future potential rewards of the LLM’s current actions.
Step 3.2: Each client linearly combines the two reward models and two Critic models according to its preferences to form a personalized reward model. This model includes immediate rewards (from the reward model1) and future potential rewards (from the Critic model1) for harmlessness, as well as immediate rewards (from the reward model2) and future potential rewards (from the Critic model2) for helpfulness.
Step 3.3: Based on the personalized reward model, each client utilizes local data to estimate the personalized loss for local LLM outputs, and uses the PPO algorithm to update the adapter and LoRA parameters of the LLM, while also updating the Critic model parameters. These adapter parameters are then transmitted over the wireless network.
Step 3.4: The server aggregates the adapter parameters from all clients to obtain a global adapter, which is then sent back to the clients.
Step 3.5: Steps 3.3-3.4 are repeated until the convergence criteria of the system are met.
@@ V Simulation Results
V-A Simulation Settings
@@
We consider a PFL system for LLMs. Suppose the system consists of ten clients and one server, where each client possesses non-iid local data and different model preferences. The server has sufficient resources. The PFL system aims to achieve personalized fine-tuning of local LLMs for all clients in wireless networks with the Rayleigh channel, adopting similar settings to those presented in [12]. Particularly, the SNR is set to 5dB, the transmission power is set to 1W, and the bandwidth is set to 1MHz.
We evaluate the proposed PWFF using the Alpaca dataset[13]. During the federated instruction tuning phase, we use GPT-2 model as the local model and exchange information between different clients using 36 adapters (bottleneck size=10). Additionally, each client designs 12-36 LoRA modules (rank size=96) based on its local resources to personalize the local model. In the federated reward learning phase, we sample instructions from the instruction dataset. We then leverage the latest GPT-4 to simulate human behavior by ranking different responses based on the helpfulness and harmlessness of their instruction-based outputs, thereby generating preference data. This preference data is then used to supervise the training process of two reward models, which are constructed using the DistilBert [14] model with 66M parameters. These reward models evaluate the helpfulness and harmlessness of the LLM responses, respectively. Finally, during the federated alignment tuning phase, we construct a personalized loss function for each client’s LLM, with the weights of helpfulness and harmlessness randomly assigned, ensuring that their sum equals 1.
V-B Evaluation Results
@@
We evaluate the performance of federated instruction tuning in PWFF using the accuracy and communication delay (i.e., communication cost divided by transmission rate) per round. Fig. 4 presents the evaluation results for PWFF and its benchmarks. In vanilla FL [1], the parameters of both 36 adapters and 36 LoRAs all need to be uploaded. Fedada [15] is federated fine-tuning method using 36 adapters, and FedLora [8] is a federated fine-tuning method that exclusively incorporates 36 LoRAs. The results indicate that PWFF achieves the highest accuracy, which highlights the effectiveness of the LoRA-based personalized structure for non-iid data. Similarly, due to the fact that PWFF only requires the transmission of a part of fine-tuned parameters (adapters), it incurs a relatively lower communication delay compared to vanilla FL.
We evaluate the performance of federated alignment tuning in PWFF using reward scores (i.e., helpful score plus harmless score) and transmission energy consumption (i.e., transmission power multiplied by transmission time). Fig. 5 presents the evaluation results for PFWW and its contenders, where SFL means a fine-tuning method that uses only a single reward model (helpfulness metric) and uploads only 36 adapter parameters. PFL, on the other hand, represents personalized fine-tuning using two reward models and uploads all adapter and LoRA parameters. Shepherd is an FL method that employs LoRA for fine-tuning [6]. We can see that PWFF enables the local model to obtain higher rewards than SFL and Shepherd with the single reward model. Moreover, compared to PFL, PFIT reduces energy consumption by 43%. Shepherd, utilizing only LoRA for fine-tuning, results in the lowest energy consumption. However, this approach also affects the LLM’s performance, resulting in the lowest reward.
@@ VI-1 Wireless Aggregation and Divergence
In PFL, the collaborative training of a shared LLM by multiple participants is challenged by signal quality fluctuations in wireless networks, potentially causing communication disruptions and data loss. These instabilities can induce model divergence, with varying participant contributions to model updates. Addressing this necessitates asynchronous model aggregation strategies and equitable client selection mechanisms to ensure the effective integration of all participant contributions.
@@ VI-2 Personalization and Overfitting
Personalization, a key objective in PFL, aims to tailor the shared LLM to each client’s specific needs. However, this could potentially induce overfitting. If personalization is too detailed, the LLM may overfit to specific client data, compromising performance on other clients or tasks. To mitigate this, appropriate regularization and control during fine-tuning are required to balance personalization and the LLM’s generalization.
@@ VI-3 Communication Efficiency and Model Accuracy
PFL’s communication and collaboration among multiple participants can lead to significant overhead, especially in scenarios with numerous participants. Frequent communication increases latency and resource consumption. Further, communication instability can cause loss or delay of model updates. Addressing this requires the design of efficient communication protocols and strategies to reduce overhead while ensuring reliable model parameter transmission.
Figure 4: Accuracy and communication delay for PWFF and its contenders.
Figure 5: Reward and energy consumption for PWFF and its contenders.
@@ VII Conclusion
In this paper, we first summarized two learning stages of LLMs in wireless networks and discussed potential solutions for combining FL with LLMs. Next, we proposed a PWFF framework for LLMs. Specifically, we used global adapters to enable information exchange between devices and incorporated local LoRAs to customize the local LLM. Then, we introduced the global partial aggregation to reduce communication delay and transmission energy consumption. Next, we designed two reward models based on helpfulness and harmlessness, and we used local multi-objective alignment to fine-tune the LLM. Finally, we carried out simulations to validate the effectiveness of the proposed PWFF framework.
@@
Feibo Jiang (jiangfb@hunnu.edu.cn) is currently an Associate Professor at Hunan Normal University, China.
Li Dong (Dlj2017@hunnu.edu.cn) is currently an Associate Professor at Hunan University of Technology and Business, China.
@@
Kezhi Wang (Kezhi.Wang@brunel.ac.uk) is a professor with the Department of Computer Science, Brunel University London, U.K.
Kun Yang (kunyang@essex.ac.uk) is currently a Chair Professor in the School of Computer Science & Electronic Engineering, University of Essex, U.K.
Cunhua Pan (cpan@seu.edu.cn) is a full professor in Southeast University, China.
Dusit Niyato (dniyato@ntu.edu.sg) is a professor in the College of Computing and Data Science, Nanyang Technological University, 639798 Singapore.